Directrix (directrix):

If the square root of a circle's area is 2012π, then how long is the radius of the circle?

OpenStudy (anonymous):

\[2012\sqrt{\pi}\]

sam (.sam.):

\[\sqrt{A}=2012\pi\] \[A=(2012\pi)^{?}\] \[\pi(r)^{2}=(2012\pi)^{?}\] \[r=\sqrt{(2012\pi)^{2}/\pi}\]

Directrix (directrix):

@.Sam --> and the final answer is ?

sam (.sam.):

3566.2

sam (.sam.):

I gtg now

Directrix (directrix):

@ Everyone -- Have the answer in exact form?

OpenStudy (anonymous):

check the first answer from ghass

Directrix (directrix):

I saw it. Interesting. Hey, how about cranking out the exact answer.

OpenStudy (anonymous):

sqrt ( π * r^2) = 2012π -> ( π * r^2) = (2012 )^2 * π^2 -> r^2 = (2012 )^2 * π => r = 2012 sqrt (π) = 6317.7